# Consistencies and inconsistencies in redshift-independent distances

### v1

@@ Abstract
Redshift-independent distances underpin much of astrophysics, and there exists a plethora of methods to estimate them. However, the extent to which the distances they imply are consistent, while crucial for the integrity of the distance ladder, has been little explored. We construct a statistical framework to assess both internal (between measurements with the same method) and external (between-method) consistency by comparing differences between distances to their quoted statistical uncertainties in the NASA/IPAC Extragalactic Database of Distances (NED-D). 66 of the 76 indicators in NED-D are amenable to a consistency test by having at least two measurements to the same galaxy or at least one measurement to a galaxy also measured by another method. We find that only 12 of these methods produce self-consistent distances across literature determinations, of which 7 are also consistent with distances to the same galaxies measured by all other methods. The most consistent 6 methods (M-stars luminosity, Novae, Masers, Globular Cluster Fundamental Plane, O- and B-type Supergiants and BL Lac Luminosity) also give similar average distances to the mean of all indicators, while the 7th (Proper Motion) underestimates distances relative to the mean by 17.1%. We also investigate consistency of Cepheid distances in the SH0ES 2022 catalogue, finding no evidence for unaccounted-for systematics. Our NED-D results imply that considerable work remains to obtain reliable distances by a multitude of methods, a crucial endeavour for constructing a multiply cross-checked and fully robust distance ladder.
@@ 1 Introduction
Assuming a cosmological model (e.g. $\Lambda$CDM) with calibrated values for its free parameters, distances in the Universe may be estimated from objects’ redshifts alone. However, within the local $\sim$100 Mpc redshifts provide inaccurate distance estimates because galaxies’ velocities are not predominantly due to the expansion of the Universe. There is a significant contribution from “peculiar velocities”, galaxies’ individual motions superimposed on the Hubble flow generated by gravitational interaction with their cosmic environment (for reviews see Strauss & Willick 1995; Turner 2024). If distances to such galaxies can be measured redshift-independently then peculiar and expansion velocities can be determined separately. The pattern of peculiar velocities provides evidence concerning the large-scale structure of the local Universe (Dupuy & Courtois, 2023; Valade et al., 2024), the growth rate of cosmic structure (Boruah et al., 2020; Said et al., 2020; Stiskalek et al., 2025), extensions to General Relativity (Lyall et al., 2023, 2024), the large-scale isotropy of the Universe (e.g. Feindt et al. 2013; Watkins et al. 2023) and the typicality or otherwise of the Milky Way’s place in the Universe (Bovy et al., 2009; McMillan, 2016).
@@ 3.1 Quantifying consistency
where $\mu_{i,j}$ are the distance moduli, and $\Sigma$ is the covariance between the two measurements. In this case $\Sigma_{ij}$ is a number given by $\sigma_{\mu_{i}}^{2}+\sigma_{\mu_{j}}^{2}-2\,\text{Cov}(\mu_{i},\mu_{j})$ where $\sigma_{\mu_{i,j}}$ are the diagonal uncertainty terms and $\text{Cov}(\mu_{i},\mu_{j})$ the cross-term between the two measurements. Note that $\text{Cov}(\mu_{i},\mu_{j})$ is assumed to be zero for all NED-D measurements, and hence only plays a role in the SH0ES analysis. Eq. 3 gives the Mahalanobis distance (Mai-Jalanobis, 1936) describing the difference in units of $\sigma$ between the measurements. It would be distributed as a standard half-normal if $\mu_{i}$ and $\mu_{j}$ were scattered from a common true value by their uncertainties (positive part of the standard normal $\mathcal{N}(\mu=0,\sigma^{2}=1)$ for $x\geq 0$). [^5] Eq. 3 may be compared to eq. 1 of Singh (2021), $\Delta_{12}=|\mu_{1}-\mu_{2}|/\sigma_{\mu_{1}}$). This has the disadvantage of not being symmetric under $\mu_{1}\leftrightarrow\mu_{2}$, and of comparing a random variable to a fixed number in its treatment of the uncertainties.
@@ 3.2 Extracting insight from inconsistency
For the case of external consistency the situation is more complex as an inconsistency could derive from inconsistent means of the $\mu_{i}$ and $\mu_{j}$ distributions rather than anything to do with the uncertainties. To disentangle the possible causes, we repeat the consistency analysis with two modifications. First, for each galaxy in the method pair, we computed the mean distance modulus over the measurements of each method. We then shift the mean of one method onto that of the other and repeat the consistency test. If this transformation produces consistency, the original inconsistency must have been due to inconsistent means. The second modification is instead of shifting the means to rescale the uncertainties by the relative standard deviations of the $\mu$ distributions. If this produces consistency, the original inconsistency must have come from the uncertainties. Note that having inconsistent means and uncertainties are not mutually exclusive. It is possible for a given method pair to have both, in which case neither of the above modifications will produce consistency by itself, but combining both of them will yield an acceptable $p$-value. If even in this case the test returns inconsistency, this must be due to other effects such as outliers, non-Gaussian uncertainties or systematic errors. We call this an “essential inconsistency,” in the sense that it cannot (easily) be removed.
Table 1: Results for the standard candle distance indicators. We present the number of galaxies having at least two measurements by a given method, the number of measurement pairs in these galaxies, the total number of measurements in NED-D along with the number of papers from which these measurements derive, the KLD $p$-value and external consistency statistics. The $p$-value is shown in green if it is greater than 0.05 (indicating consistency), orange if it is between 0 and 0.05 (indicating inconsistency), and red if it is 0 (indicating severe inconsistency). In cases of inconsistency, we indicate the $p$-value after rescaling the uncertainties to a standard half-normal in parenthesis. If this produces consistency, we indicate whether the uncertainties were originally underestimated (downward-pointing arrow) or overestimated (upward-pointing arrow). The final column shows the fraction of other methods with which the method in question is externally consistent, either out-of-the-box, after making the means consistent of after making the uncertainties consistent or both. Methods that are internally consistent and members of the maximal externally consistent set are shown in bold. (See text for full details.)
@@
BL Lac Luminosity 2 4 20 / 11 0.9705 1 - 1 - 1 - 1 / 1
RSV Stars 2 4 9 / 6 0.8298 9 - 22 - 27 - 27 / 28
OB Stars 2 2 4 / 3 0.7383 27 - 30 - 30 - 31 / 31
Novae 6 14 18 / 11 0.4382 35 - 40 - 43 - 43 / 46
Miras 6 119 36 / 27 0.1121 20 - 24 - 32 - 32 / 39
RV Stars 1 10 5 / 1 0.0845 6 - 7 - 14 - 14 / 18
Statistical 14 168 292 / 27 0.0795 15 - 22 - 35 - 35 / 54
Brightest Stars 21 127 129 / 41 0.074 21 - 26 - 40 - 40 / 44
M Stars 1 3 6 / 5 0.0686 18 - 24 - 32 - 32 / 33
Delta Scuti 2 29 12 / 2 0.0395 (0.8241) $\uparrow$ 14 - 16 - 27 - 27 / 31
Horizontal Branch 22 139 110 / 53 0.0037 (0.1495) $\downarrow$ 19 - 25 - 36 - 36 / 40
PNLF 45 285 183 / 57 0.0016 (0.0019) 23 - 26 - 41 - 41 / 52
AGN time lag 11 13 32 / 3 0.0 (0.7099) $\downarrow$ 2 - 3 - 6 - 6 / 6
Carbon Stars 7 182 82 / 13 0.0 (0.4555) $\uparrow$ 25 - 27 - 33 - 33 / 38
Cepheids 115 69496 10956 / 325 0.0 (0.0) 8 - 12 - 26 - 26 / 47
CMD 123 41733 1265 / 170 0.0 (0.0) 15 - 23 - 34 - 35 / 51
@@
GCLF 176 2288 781 / 96 0.0 (0.0) 15 - 21 - 31 - 31 / 40
RR Lyrae 2443 16262 26409 / 216 0.0 (0.0) 16 - 23 - 31 - 32 / 43
Red Clump 24 3420 236 / 63 0.0 (0.0) 16 - 19 - 28 - 28 / 40
SNIa SDSS 1849 7687 6054 / 1 0.0 (0.0) 0 - 0 - 1 - 1 / 4
SZ effect 78 426 283 / 19 0.0 (0.0) 6 - 12 - 15 - 15 / 16
SBF 406 4403 1781 / 67 0.0 (0.0) 19 - 23 - 31 - 31 / 42
TRGB 307 8537 1736 / 336 0.0 (0.0) 15 - 26 - 34 - 35 / 54
Type II Cepheids 7 435 62 / 22 0.0 (0.3613) $\downarrow$ 18 - 24 - 40 - 40 / 44
SNIa 4978 129581 30843 / 140 0.0 (0.0) 10 - 14 - 22 - 22 / 38
AGB 0 0 2 / 2 - 22 - 23 - 23 - 23 / 23
B Stars 0 0 2 / 2 - 28 - 28 - 28 - 28 / 30
Black Hole 0 0 17 / 2 - 4 - 6 - 10 - 11 / 11
Blue Supergiant 0 0 2 / 1 - 11 - 12 - 25 - 25 / 25
BCG 0 0 239 / 1 - 9 - 9 - 11 - 11 / 11
FGLR 0 0 7 / 7 - 9 - 15 - 21 - 21 / 21
GC SBF 0 0 3 / 1 - 13 - 14 - 18 - 18 / 19
HII LF 0 0 17 / 2 - 6 - 11 - 12 - 12 / 15
SX Phe Stars 0 0 3 / 2 - 10 - 11 - 13 - 13 / 13
Subdwarf fitting 0 0 1 / 1 - 19 - 22 - 22 - 22 / 22
SNII radio 0 0 24 / 2 - 3 - 9 - 15 - 16 / 17
White Dwarfs 0 0 1 / 1 - 21 - 22 - 21 - 22 / 22
Wolf-Rayet 0 0 1 / 1 - 0 - 3 - 5 - 5 / 5
Grav. Wave 0 0 8 / 8 - 4 - 4 - 4 - 4 / 4
@@ B Stars 0 0 2 / 2 - 28 - 28 - 28 - 28 / 30
Proper Motion 1 6 5 / 3 0.8918 24 - 26 - 29 - 30 / 32
Maser 6 29 24 / 20 0.6253 28 - 32 - 35 - 36 / 36
Dwarf Galaxy Diameter 8 8 16 / 1 0.0004 (0.582) $\uparrow$ 17 - 17 - 21 - 21 / 21
G Lens 15 70 86 / 14 0.0002 (0.6975) $\downarrow$ 1 - 1 - 3 - 3 / 3
Eclipsing Binary 49 3397 273 / 48 0.0 (0.0007) 22 - 23 - 34 - 34 / 42
HII region diameter 19 25 53 / 7 0.0 (0.0261) 7 - 20 - 42 - 42 / 42
Ring Diameter 47 47 212 / 1 0.0 (0.2025) $\uparrow$ 10 - 11 - 18 - 18 / 18
SNII optical 309 6090 1622 / 89 0.0 (0.0) 17 - 22 - 32 - 32 / 42
GC radius 0 0 101 / 8 - 22 - 25 - 34 - 34 / 37
Grav. Stability Gas. Disk 0 0 2 / 1 - 1 - 3 - 4 - 4 / 4
Orbital Mech. 0 0 3 / 3 - 20 - 22 - 22 - 22 / 22
GC FP 1 3 3 / 3 0.2469 16 - 17 - 18 - 19 / 19
GeV TeV ratio 4 8 30 / 11 0.0385 (0.7002) $\downarrow$ 1 - 1 - 1 - 1 / 1
Faber-Jackson 427 1947 1424 / 6 0.0101 (0.6445) $\uparrow$ 5 - 10 - 25 - 25 / 29
Diameter 3 3 6 / 2 0.0089 (0.0504) $\uparrow$ 18 - 20 - 26 - 26 / 31
D-Sigma 551 2901 1995 / 9 0.0 (0.0) 9 - 12 - 14 - 14 / 25
FP 967 1675 130214 / 22 0.0 (0.0) 10 - 13 - 19 - 19 / 28
IRAS 405 640 2947 / 2 0.0 (0.0) 4 - 5 - 20 - 21 / 35
Mass Model 2 4 7 / 7 0.0 (0.8942) $\downarrow$ 0 - 1 - 10 - 10 / 10
Tully-Fisher 9107 231663 56436 / 78 0.0 (0.0) 13 - 24 - 31 - 32 / 48
H I + optical distribution 0 0 1 / 1 - 0 - 1 - 1 - 1 / 1
LSB galaxies 0 0 2 / 1 - 7 - 8 - 9 - 9 / 10
Magnitude 0 0 111 / 3 - 13 - 31 - 38 - 38 / 40
Sosies 0 0 287 / 3 - 10 - 12 - 20 - 20 / 21
Tertiary 0 0 284 / 4 - 6 - 11 - 14 - 14 / 22
Tully est 0 0 1434 / 1 - 17 - 17 - 28 - 28 / 41

### v2

@@ Abstract
Redshift-independent distances underpin much of astrophysics, and there exists a plethora of methods to estimate them. However, the extent to which the distances they imply are consistent, while crucial for the integrity of the distance ladder, has been little explored. We construct a statistical framework to assess both internal (between measurements with the same method) and external (between-method) consistency by comparing differences between distances to their quoted statistical uncertainties in the NASA/IPAC Extragalactic Database of Distances (NED-D). 66 of the 76 indicators in NED-D are amenable to a consistency test by having at least two measurements to the same galaxy or at least one measurement to a galaxy also measured by another method. We find that only 12 of these methods produce self-consistent distances across literature determinations, of which 7 are also consistent with distances to the same galaxies measured by all other methods. The most consistent 6 methods (M-stars luminosity, Novae, Masers, Globular Cluster Fundamental Plane, O- and B-type Supergiants and BL Lac Luminosity) also give similar average distances to the mean of all indicators, while the 7th (Proper Motion) underestimates distances relative to the mean by 17.1%. We also investigate consistency of Cepheid distances in the SH0ES 2022 catalogue, finding no evidence for unaccounted-for systematics. Our NED-D results imply that considerable work remains to obtain reliable distances by a multitude of methods, a crucial endeavour for constructing a multiply cross-checked and fully robust distance ladder. We make our code publicly available.
@@ 1 Introduction
Assuming a cosmological model (e.g. $\Lambda$CDM) with calibrated values for its free parameters, distances in the Universe may be estimated from objects’ redshifts alone. However, within the local $\sim$100 Mpc redshifts provide inaccurate distance estimates because galaxies’ velocities are not predominantly due to the expansion of the Universe. There is a significant contribution from “peculiar velocities”, galaxies’ individual motions superimposed on the Hubble flow generated by gravitational interaction with their cosmic environment (for reviews see Strauss & Willick 1995; Turner 2024). If distances to such galaxies can be measured redshift-independently then peculiar and expansion velocities can be determined separately. The pattern of peculiar velocities provides evidence concerning the large-scale structure of the local Universe (Dupuy & Courtois, 2023; Valade et al., 2024), the growth rate of cosmic structure (Boruah et al., 2020; Said et al., 2020; Stiskalek et al., 2025), extensions to General Relativity (Lyall et al., 2023, 2024), the large-scale isotropy of the Universe (e.g. Feindt et al. 2013; Watkins et al. 2023) and the typicality or otherwise of the Milky Way’s place in the Universe (Bovy et al., 2009; McMillan, 2016). It also allows the distance–redshift relation to test, rather than assume, cosmological models.
@@ 3.1 Quantifying consistency
where $\mu_{i,j}$ are the distance moduli, and $\Sigma$ is the covariance between the two measurements. In this case $\Sigma_{ij}$ is a number given by $\sigma_{\mu_{i}}^{2}+\sigma_{\mu_{j}}^{2}-2\,\text{Cov}(\mu_{i},\mu_{j})$ where $\sigma_{\mu_{i,j}}$ are the diagonal uncertainty terms and $\text{Cov}(\mu_{i},\mu_{j})$ the cross-term between the two measurements. Note that $\text{Cov}(\mu_{i},\mu_{j})$ is assumed to be zero for all NED-D measurements, and hence only plays a role in the SH0ES analysis. Eq. 3 gives the Mahalanobis distance (Mai-Jalanobis, 1936) describing the difference in units of $\sigma$ between the measurements. It would be distributed as a standard half-normal if $\mu_{i}$ and $\mu_{j}$ were scattered from a common true value by their uncertainties (positive part of the standard normal $\mathcal{N}(\mu=0,\sigma^{2}=1)$ for $x\geq 0$). [^5] Eq. 3 may be compared to eq. 1 of Singh (2021), $\Delta_{12}=|\mu_{1}-\mu_{2}|/\sigma_{\mu_{1}}$). This is not symmetric under $\mu_{1}\leftrightarrow\mu_{2}$, although the combination effectively used in that work, $(\Delta_{12}+\Delta_{21})/2$, is. Our formula has the justification of being derived from the convolution of two Gaussians assumed to be independent.
@@
Our choice to define the null hypothesis as consistency between the measurements means that significant evidence is required before concluding inconsistency. One could imagine instead defining the null hypothesis as inconsistency, so that significant evidence would be required to conclude consistency. Our choice has two advantages. First, while “consistency” is a well-defined model (each galaxy has an underlying true distance, which produces the various measurements through scattering by the observational uncertainties), inconsistency is not: there is an infinity of ways in which two sets of measurements may be inconsistent. It is therefore difficult to see how a null hypothesis of inconsistency could be unambiguously defined. Second, our choice makes our results conservative in the sense that the inconsistencies that we identify are significant by construction. We caution however that while we apply the term “consistency” to any comparison yielding $p>0.05$, we really mean only that the null hypothesis of consistency cannot be robustly ruled out. A $p$-value greater than 0.05 may still favour inconsistency. In some sense our results therefore provide a lower bound on the degree of inconsistency in the datasets we analyse.
@@ 3.2 Extracting insight from inconsistency
For the case of external consistency the situation is more complex as an inconsistency could derive from inconsistent means of the $\mu_{i}$ and $\mu_{j}$ distributions rather than anything to do with the uncertainties. To disentangle the possible causes, we repeat the consistency analysis with two modifications. First, for each galaxy in the method pair, we computed the mean distance modulus over the measurements of each method. We then shift the mean of one method onto that of the other and repeat the consistency test. If this transformation produces consistency, the original inconsistency must have been due to inconsistent means. The second modification is instead of shifting the means to rescale the uncertainties by the relative standard deviations of the $\mu$ distributions. If this produces consistency, the original inconsistency must have come from the uncertainties. Note that having inconsistent means and uncertainties are not mutually exclusive. In some cases consistency may be achieved either by shifting the means or rescaling the uncertainties; in others doing both is required. If the test returns inconsistency even after both of these are done, it must be due to other effects such as outliers, non-Gaussian uncertainties or systematic errors. We call this an “essential inconsistency,” indicating that it cannot (easily) be removed.
Table 1: Results for the standard candle distance indicators. We present the number of galaxies having at least two measurements by a given method, the number of measurement pairs in these galaxies, the total number of measurements in NED-D along with the number of papers from which these measurements derive, the KLD $p$-value and external consistency statistics. The $p$-value is shown in green if it is greater than 0.05 (indicating consistency, i.e. not enough evidence to reject the null hypothesis), orange if it is between 0 and 0.05 (indicating inconsistency), and red if it is 0 (indicating severe inconsistency). In cases of inconsistency, we indicate the $p$-value after rescaling the uncertainties to a standard half-normal in parenthesis. If this produces consistency, we indicate whether the uncertainties were originally underestimated (downward-pointing arrow) or overestimated (upward-pointing arrow). The final column shows the fraction of other methods with which the method in question is externally consistent: the first number gives the number of cases that are consistent out-of-the-box, the second the number that are made consistent by shifting the means but not by rescaling the uncertainties, the third the number that are made consistent by rescaling the uncertainties but not by shifting the means, the fourth the number that are made consistent by either shifting the means or rescaling the uncertainties, and the fifth the number that are made consistent only by both shifting the means and rescaling the uncertainties. The final number shows the total number of other methods with which a given method is compared, such that this number minus the sum of the others is the number of essential inconsistencies.
Methods that are internally consistent and members of the maximal externally consistent set are shown in bold. (See text for full details.)
@@
BL Lac Luminosity 2 4 20 / 11 0.9705 1 - 0 - 0 - 0 - 0 / 1
RSV Stars 2 4 9 / 6 0.8298 9 - 0 - 3 - 4 - 0 / 28
OB Stars 2 2 4 / 3 0.7383 27 - 0 - 1 - 0 - 0 / 31
Novae 6 14 18 / 11 0.4382 35 - 0 - 0 - 3 - 1 / 46
Miras 6 119 36 / 27 0.1121 20 - 0 - 4 - 2 - 0 / 39
RV Stars 1 10 5 / 1 0.0845 6 - 0 - 1 - 0 - 0 / 18
Statistical 14 168 292 / 27 0.0795 15 - 0 - 10 - 3 - 0 / 54
Brightest Stars 21 127 129 / 41 0.074 21 - 0 - 6 - 3 - 2 / 44
M Stars 1 3 6 / 5 0.0686 18 - 0 - 1 - 1 - 0 / 33
Delta Scuti 2 29 12 / 2 0.0395 (0.8241) $\uparrow$ 14 - 0 - 4 - 1 - 0 / 31
Horizontal Branch 22 139 110 / 53 0.0037 (0.1495) $\downarrow$ 19 - 0 - 4 - 4 - 0 / 40
PNLF 45 285 183 / 57 0.0016 (0.0019) 23 - 0 - 12 - 3 - 1 / 52
AGN time lag 11 13 32 / 3 0.0 (0.7099) $\downarrow$ 2 - 0 - 3 - 1 - 0 / 6
Carbon Stars 7 182 82 / 13 0.0 (0.4555) $\uparrow$ 25 - 0 - 1 - 1 - 0 / 38
Cepheids 115 69496 10956 / 325 0.0 (0.0) 8 - 0 - 6 - 2 - 1 / 47
CMD 123 41733 1265 / 170 0.0 (0.0) 15 - 0 - 6 - 5 - 1 / 51
@@
GCLF 176 2288 781 / 96 0.0 (0.0) 15 - 0 - 4 - 1 - 0 / 40
RR Lyrae 2443 16262 26409 / 216 0.0 (0.0) 16 - 1 - 4 - 2 - 2 / 43
Red Clump 24 3420 236 / 63 0.0 (0.0) 16 - 0 - 4 - 1 - 1 / 40
SNIa SDSS 1849 7687 6054 / 1 0.0 (0.0) 0 - 0 - 2 - 0 - 0 / 4
SZ effect 78 426 283 / 19 0.0 (0.0) 6 - 0 - 0 - 1 - 0 / 16
SBF 406 4403 1781 / 67 0.0 (0.0) 19 - 0 - 8 - 2 - 1 / 42
TRGB 307 8537 1736 / 336 0.0 (0.0) 15 - 1 - 3 - 9 - 0 / 54
Type II Cepheids 7 435 62 / 22 0.0 (0.3613) $\downarrow$ 18 - 1 - 6 - 1 - 0 / 44
SNIa 4978 129581 30843 / 140 0.0 (0.0) 10 - 0 - 5 - 2 - 0 / 38
AGB 0 0 2 / 2 - 22 - 0 - 0 - 0 - 0 / 23
B Stars 0 0 2 / 2 - 28 - 0 - 0 - 0 - 0 / 30
Black Hole 0 0 17 / 2 - 4 - 1 - 2 - 1 - 0 / 11
Blue Supergiant 0 0 2 / 1 - 11 - 0 - 0 - 0 - 0 / 25
BCG 0 0 239 / 1 - 9 - 0 - 2 - 0 - 0 / 11
FGLR 0 0 7 / 7 - 9 - 0 - 3 - 4 - 0 / 21
GC SBF 0 0 3 / 1 - 13 - 0 - 2 - 0 - 0 / 19
HII LF 0 0 17 / 2 - 6 - 0 - 0 - 1 - 0 / 15
SX Phe Stars 0 0 3 / 2 - 10 - 0 - 0 - 0 - 0 / 13
Subdwarf fitting 0 0 1 / 1 - 19 - 0 - 0 - 0 - 0 / 22
SNII radio 0 0 24 / 2 - 3 - 1 - 6 - 4 - 0 / 17
White Dwarfs 0 0 1 / 1 - 21 - 0 - 0 - 0 - 0 / 22
Wolf-Rayet 0 0 1 / 1 - 0 - 0 - 1 - 3 - 0 / 5
Grav. Wave 0 0 8 / 8 - 4 - 0 - 0 - 0 - 0 / 4
@@ B Stars 0 0 2 / 2 - 28 - 0 - 0 - 0 - 0 / 30
Proper Motion 1 6 5 / 3 0.8918 24 - 2 - 3 - 0 - 0 / 32
Maser 6 29 24 / 20 0.6253 28 - 0 - 4 - 2 - 0 / 36
Dwarf Galaxy Diameter 8 8 16 / 1 0.0004 (0.582) $\uparrow$ 17 - 0 - 3 - 0 - 0 / 21
G Lens 15 70 86 / 14 0.0002 (0.6975) $\downarrow$ 1 - 0 - 2 - 0 - 0 / 3
Eclipsing Binary 49 3397 273 / 48 0.0 (0.0007) 22 - 0 - 2 - 1 - 0 / 42
HII region diameter 19 25 53 / 7 0.0 (0.0261) 7 - 0 - 15 - 7 - 0 / 42
Ring Diameter 47 47 212 / 1 0.0 (0.2025) $\uparrow$ 10 - 0 - 6 - 1 - 0 / 18
SNII optical 309 6090 1622 / 89 0.0 (0.0) 17 - 0 - 4 - 3 - 0 / 42
GC radius 0 0 101 / 8 - 22 - 0 - 9 - 2 - 0 / 37
Grav. Stability Gas. Disk 0 0 2 / 1 - 1 - 0 - 0 - 0 - 0 / 4
Orbital Mech. 0 0 3 / 3 - 20 - 0 - 0 - 2 - 0 / 22
GC FP 1 3 3 / 3 0.2469 16 - 1 - 0 - 0 - 0 / 19
GeV TeV ratio 4 8 30 / 11 0.0385 (0.7002) $\downarrow$ 1 - 0 - 0 - 0 - 0 / 1
Faber-Jackson 427 1947 1424 / 6 0.0101 (0.6445) $\uparrow$ 5 - 0 - 11 - 5 - 0 / 29
Diameter 3 3 6 / 2 0.0089 (0.0504) $\uparrow$ 18 - 0 - 2 - 0 - 0 / 31
D-Sigma 551 2901 1995 / 9 0.0 (0.0) 9 - 0 - 4 - 1 - 0 / 25
FP 967 1675 130214 / 22 0.0 (0.0) 10 - 0 - 5 - 2 - 0 / 28
IRAS 405 640 2947 / 2 0.0 (0.0) 4 - 1 - 16 - 0 - 0 / 35
Mass Model 2 4 7 / 7 0.0 (0.8942) $\downarrow$ 0 - 0 - 0 - 0 - 0 / 10
Tully-Fisher 9107 231663 56436 / 78 0.0 (0.0) 13 - 1 - 5 - 3 - 0 / 48
H I + optical distribution 0 0 1 / 1 - 0 - 0 - 0 - 0 - 0 / 1
LSB galaxies 0 0 2 / 1 - 7 - 0 - 0 - 1 - 0 / 10
Magnitude 0 0 111 / 3 - 13 - 0 - 5 - 12 - 0 / 40
Sosies 0 0 287 / 3 - 10 - 0 - 7 - 1 - 0 / 21
Tertiary 0 0 284 / 4 - 6 - 0 - 3 - 2 - 0 / 22
Tully est 0 0 1434 / 1 - 17 - 0 - 8 - 0 - 0 / 41